I have a question that I think I know the answer to but I need someone to look it over and tell me if it is right. Solve the following equation: 6/11 =114/n.
I think the answer is n=684? Can someone tell me if that was right?

Answer by sabanasir(22)

listen:
6/11 =114/n
ok now see this: when u cross multiply u'll get:
6 * n = 11 * 114
now: leaving 'n' alone by multiplying 1/6 on both sides u get:
6* 1/6 * n =114 * 11 * 1/6
=> n = (114 * 11) / 6
=> n= 209
the answser u provided is wrong